The shortest side of a triangle is one half the length of the middle side, plus 2 cm. The longest side of the triangle is one and one-half times the length of the middle side, less 2 cm. The perimeter of the triangle is 60 cm. How long are each of the sides?
Answer by drk(1908)

S = (1/2)m + 2
m = m
L = (3/2)m - 2
P = 60
----
P = short + middle + long
60 = (1/2)m + 2 + m +(3/2)m - 2
combine like terms to get
60 = 3m
divide to get
m = 20
s = 12
L = 28